P(sapphire wins) = .22
P(sapphire loses) = 1-.22 = .78
P( emerald wins) = .47
P( emerald loses) = 1- .47 = .53
P ( sapphire loses & emerald loses ) = P( sapphire loses) P(emerald loses) = 0.780.53 = 0.4134
Which multiplied by 100 means 41.34%. I hope that this is the answer that you were looking for and it has helped you.